OpenStudy (anonymous):

Hey! xD (2x)^3 How do you solve when x=6?

OpenStudy (anonymous):

how much is 2*x when x=6 ?

OpenStudy (anonymous):

just plug 6 in place of x

OpenStudy (anonymous):

12^3

OpenStudy (anonymous):

I know the plug in part. I just don't speak exponent....

OpenStudy (anonymous):

12^3=12*12*12

OpenStudy (anonymous):

12^3

OpenStudy (anonymous):

12*12*12=1728

OpenStudy (anonymous):

Ok, fiddle......so 1728? Is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

12^1=12 12^2=12*12 12^3=12*12*12

OpenStudy (anonymous):

You don't simplify?

OpenStudy (anonymous):

nothing to simplify further in 1728

OpenStudy (anonymous):

Ok, on to my next problem!!!!! I post on left. ;D